Period of [sinx]

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
f(x) = [sin(x)]
f(x) = 0, 0\leq x< \frac{\pi }{2}
=1, x = \frac{\pi }{2}
= 0, \frac{\pi }{2}< x\leq \pi
=-1, \pi < x \leq 2\pi
Same pattern will repeat after this.
Time period of f(x)
2\pi
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty